OpenStudy (aaronandyson):

A two digit positive number such that the products of its numbers is 6.If 9 added the number the digits, interchange places.Find the original number.

OpenStudy (aaronandyson):

@Data_LG2

OpenStudy (aaronandyson):

@The_Joker<3

OpenStudy (anonymous):

Let \(10a+b\) denote the original number, with \(a\) one of the numerals 1-9 and \(b\) one of \(0-9\). You have the following system of equations: \[\begin{cases}ab=6\\ 10a+b+9=10b+a\end{cases}\]

OpenStudy (anonymous):

Oh, and now you know \(b\) must be a numeral from 1-9, it can no longer be 0.

OpenStudy (anonymous):

{ a b = 6, 9 + a + b = 10 b + a} a = 6, b = 1